OpenStudy (anonymous):

What is the range of y = -3x2 + 2?

OpenStudy (acxbox22):

in set notation?

OpenStudy (anonymous):

Well first, do you know what the difference between range and domain are @JBM?

OpenStudy (anonymous):

And second, are you saying \[y = 3x ^{2} + 2\] or y = 3x2 + 2?

OpenStudy (anonymous):

Is it squared or not? The operation on x.

OpenStudy (anonymous):

I'm guessing it is squared, which yields a graph like this,

OpenStudy (anonymous):

Range refers to all of the y-values that validate your function.

OpenStudy (anonymous):

If you want me to continue, please say so.

OpenStudy (triciaal):

when x = 0 y = 2 when y = 0 then -3x^2 +2 =0 x^2 = -2/-3 so x = =/-rt 2/3 plotting some point when x = -2 y =-10 when x = -1 y = -1 when x = 1 y = -3 +2 = -1 when x = 2 y = -10  for x -inf to -rt2/3 y <0 for x -rt 2/3 to + rt 2/3 0<y <2 for x >rt 2/3 y <0 Range is the set of y values - inf to 2